40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 241     Entry time: Wed Jan 16 14:09:45 2008
Author: rob 
Type: Update 
Category: LSC 
Subject: tuesday's locking 

I got a little further with the locking (PRFPMI) last night, after discovering that the cable going from the CM board to the MC board was unplugged at the MC side. This explains why we weren't able to engage the AO path last night. Tonight, I got up to the point where DARM is handed off to OMC transmission, a step which repeatedly failed.
Eventually I realized that although all the lights are the green, the OMC Trans signal was not being updated in the LSC's memory. I suspect this is because the c1ass machine was powered down. Work continues.
ELOG V3.1.3-